OpenStudy (anonymous):

Solve for x: 1/4(2x - 14) = 4

pooja195 (pooja195):

distribute the 1/4 1/4(2x - 14)

OpenStudy (anonymous):

1/4(2x)-1/4(14)

pooja195 (pooja195):

yes can you solve those?

pooja195 (pooja195):

by multiplying?

pooja195 (pooja195):

1/2x-7/2= 4 here i did that step for you :) since we are isolating we add 7/2 to both sides 1/2x=4+7/2 so what does 4+7/2 = ?

OpenStudy (anonymous):

5.5?

pooja195 (pooja195):

Not quite 1/2x=15/2 it would be 15/2 <---use a calculator to help you :) and that as a decimal is 7.5 whats 15/2 divde by 1/2 ?

OpenStudy (anonymous):

3.75

OpenStudy (pratyush5):